# DUSt3R: Geometric 3D Vision Made Easy

### v1

@@ 1 Introduction
Unconstrained image-based dense 3D reconstruction from multiple views is one of a few long-researched end-goals of computer vision [24, 89, 71]. In a nutshell, the task aims at estimating the 3D geometry and camera parameters of a particular scene, given a set of photographs of this scene. Not only does it have numerous applications like mapping [72, 13] , navigation [15], archaeology [86, 132], cultural heritage preservation [38], robotics [78], but perhaps more importantly, it holds a fundamentally special place among all 3D vision tasks. Indeed, it subsumes nearly all of the other geometric 3D vision tasks. Thus, modern approaches for 3D reconstruction consists in assembling the fruits of decades of advances in various sub-fields such as keypoint detection [62, 96, 26, 28] and matching [99, 119, 10, 59], robust estimation [3, 10, 180], Structure-from-Motion (SfM) and Bundle Adjustment (BA) [20, 105, 58], dense Multi-View Stereo (MVS) [106, 175, 157, 138], etc.
@@ 2 Related Work
MultiView Stereo (MVS) is the task of densely reconstructing visible surfaces, which is achieved via triangulation between multiple viewpoints. In the classical formulation of MVS, all camera parameters are supposed to be provided as inputs. The fully handcrafted [32, 174, 34, 106, 146], the more recent scene optimization based [75, 76, 142, 31, 162, 70, 145, 147], or learning based [52, 160, 179, 163, 85, 64] approaches all depend on camera parameter estimates obtained via complex calibration procedures, either during the data acquisition [1, 108, 23, 165] or using Structure-from-Motion approaches [105, 47] for in-the-wild reconstructions. Yet, in real-life scenarios, the inaccuracy of pre-estimated camera parameters can be detrimental for these algorithms to work properly. In this work, we propose instead to directly predict the geometry of visible surfaces without any explicit knowledge of the camera parameters.
@@ 3 Method
$X^{n,m}=P_{m}P_{n}^{-1}h\left(X^{n}\right)\vskip-5.69054pt$ (1)
@@ 3.1 Overview
$F^{1}=\text{Encoder}(I^{1}),F^{2}=\text{Encoder}(I^{2}).\vskip-5.69054pt$
@@
$G_{i}^{2}=\text{DecoderBlock}^{2}_{i}\left(G_{i-1}^{2},G_{i-1}^{1}\right),\vskip-8.53581pt$
@@ 3.2 Training Objective
$\ell_{\text{regr}}(v,i)=\left\|\frac{1}{z}X^{v,1}_{i}-\frac{1}{\bar{z}}\bar{X}^{v,1}_{i}\right\|.\vskip-2.84526pt$ (2)
@@
$\text{norm}(X^{1},X^{2})=\frac{1}{|\mathcal{D}^{1}|+|\mathcal{D}^{2}|}\sum_{v\in\{1,2\}}\,\sum_{i\in\mathcal{D}^{v}}\left\|X^{v}_{i}\right\|.\vskip-5.69054pt$ (3)
@@
$\mathcal{L}_{\text{conf}}=\sum_{v\in\{1,2\}}\,\sum_{i\in\mathcal{D}^{v}}C^{v,1}_{i}\ell_{\text{regr}}(v,i)-\alpha\log C^{v,1}_{i},\vskip-5.69054pt$ (4)
@@ 3.3 Downstream Applications
$f_{1}^{*}=\operatorname*{arg\,min}_{f_{1}}\sum_{i=0}^{W}\sum_{j=0}^{H}C^{1,1}_{i,j}\left\|(i^{\prime},j^{\prime})-f_{1}\frac{(X^{1,1}_{i,j,0},X^{1,1}_{i,j,1})}{X^{1,1}_{i,j,2}}\right\|,\vskip-2.84526pt$
@@
$R^{*},t^{*}=\operatorname*{arg\,min}_{\sigma,R,t}\sum_{i}C^{1,1}_{i}C^{1,2}_{i}\left\|\sigma(RX^{1,1}_{i}+t)-X^{1,2}_{i}\right\|^{2},\vskip-5.69054pt$
@@ 3.4 Global Alignment
$\chi^{*}=\operatorname*{arg\,min}_{\chi,P,\sigma}\sum_{e\in\mathcal{E}}\sum_{v\in e}\sum_{i=1}^{HW}C^{v,e}_{i}\left\|\chi_{i}^{v}-\sigma_{e}P_{e}X^{v,e}_{i}\right\|.\vskip-8.53581pt$ (5)
@@ 4 Experiments with DUSt3R
Training details. During each epoch, we randomly sample an equal number of pairs from each dataset to equalize disparities in dataset sizes. We wish to feed relatively high-resolution images to our network, say 512 pixels in the largest dimension. To mitigate the high cost associated with such input, we train our network sequentially, first on $\times224$ images and then on larger 512-pixel images. We randomly select the image aspect ratios for each batch (e.g. 16/9, 4/3, etc), so that at test time our network is familiar with different image shapes. We simply crop images to the desired aspect-ratio, and resize so that the largest dimension is 512 pixels.
@@ 4.2 Multi-view Pose Estimation
Baselines and metrics. We compare DUSt3R pose estimation results, obtained either from PnP-RANSAC or global alignment, against the learning-based RelPose [176], PoseReg [139] and PoseDiffusion [139], and structure-based PixSFM [58], COLMAP+SPSG (COLMAP [106] extended with SuperPoint [26] and SuperGlue [99]). Similar to [139], we report the Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) for each image pair to evaluate the relative pose error and select a threshold $\tau=15$ to report RTA@$15$ and RRA@15@15. Additionally, we calculate the mean Average Accuracy (mAA)@30@30, defined as the area under the curve accuracy of the angular differences at $min({\rm RRA}@30,{\rm RTA}@30)$.
@@
Monodepth2 [37] $S$ 23.91 75.22 11.42 86.90 56.49 35.18 16.19 74.50 31.20 47.42
SC-SfM-Learners [6] $S$ 16.92 77.28 11.83 86.61 21.11 71.40 13.79 79.57 22.29 64.30
SC-DepthV3 [120] $S$ 14.20 81.27 11.79 86.39 12.58 88.92 12.34 84.80 16.28 79.67
MonoViT[181] $S$ - - 09.92 90.01 - - - - -
@@
Table 2: Left: Monocular depth estimation on multiple benchmarks. D-Supervised, $S$-Self-supervised, T-transfer (zero-shot). (Parentheses) refers to training on the same set. Right: Multi-view pose regression on the CO3Dv2 [93] and RealEst10K [185] with 10 random frames.
@@
[…]
@@ 4.6 Ablations
Table 4: MVS results on the DTU dataset, in $m$. Traditional handcrafted methods (a) have been overcome by learning-based approaches (b) that train on this specific domain.

### v2

@@ 1 Introduction
Unconstrained image-based dense 3D reconstruction from multiple views is one of a few long-researched end-goals of computer vision [24, 90, 72]. In a nutshell, the task aims at estimating the 3D geometry and camera parameters of a particular scene, given a set of photographs of this scene. Not only does it have numerous applications like mapping [73, 13] , navigation [15], archaeology [87, 133], cultural heritage preservation [38], robotics [79], but perhaps more importantly, it holds a fundamentally special place among all 3D vision tasks. Indeed, it subsumes nearly all of the other geometric 3D vision tasks. Thus, modern approaches for 3D reconstruction consists in assembling the fruits of decades of advances in various sub-fields such as keypoint detection [63, 97, 26, 28] and matching [100, 120, 10, 60], robust estimation [3, 10, 181], Structure-from-Motion (SfM) and Bundle Adjustment (BA) [20, 106, 59], dense Multi-View Stereo (MVS) [107, 176, 158, 139, 40], etc.
@@ 2 Related Work
MultiView Stereo (MVS) is the task of densely reconstructing visible surfaces, which is achieved via triangulation between multiple viewpoints. In the classical formulation of MVS, all camera parameters are supposed to be provided as inputs. The fully handcrafted [32, 175, 34, 107, 147], the more recent scene optimization based [76, 77, 143, 31, 163, 71, 146, 148], or learning based [53, 161, 180, 164, 86, 65, 40] approaches all depend on camera parameter estimates obtained via complex calibration procedures, either during the data acquisition [1, 109, 23, 166] or using Structure-from-Motion approaches [106, 48] for in-the-wild reconstructions. Yet, in real-life scenarios, the inaccuracy of pre-estimated camera parameters can be detrimental for these algorithms to work properly. In this work, we propose instead to directly predict the geometry of visible surfaces without any explicit knowledge of the camera parameters.
@@ 3 Method
$X^{n,m}=P_{m}P_{n}^{-1}h\left(X^{n}\right)\vspace{-2mm}$ (1)
@@ 3.1 Overview
$F^{1}=\text{Encoder}(I^{1}),F^{2}=\text{Encoder}(I^{2}).\vspace{-2mm}$
@@
$G_{i}^{2}=\text{DecoderBlock}^{2}_{i}\left(G_{i-1}^{2},G_{i-1}^{1}\right),\vspace{-3mm}$
@@ 3.2 Training Objective
$\ell_{\text{regr}}(v,i)=\left\|\frac{1}{z}X^{v,1}_{i}-\frac{1}{\bar{z}}\bar{X}^{v,1}_{i}\right\|.\vspace{-1mm}$ (2)
@@
$\text{norm}(X^{1},X^{2})=\frac{1}{|\mathcal{D}^{1}|+|\mathcal{D}^{2}|}\sum_{v\in\{1,2\}}\,\sum_{i\in\mathcal{D}^{v}}\left\|X^{v}_{i}\right\|.\vspace{-2mm}$ (3)
@@
$\mathcal{L}_{\text{conf}}=\sum_{v\in\{1,2\}}\,\sum_{i\in\mathcal{D}^{v}}C^{v,1}_{i}\ell_{\text{regr}}(v,i)-\alpha\log C^{v,1}_{i},\vspace{-2mm}$ (4)
@@ 3.3 Downstream Applications
$f_{1}^{*}=\operatorname*{arg\,min}_{f_{1}}\sum_{i=0}^{W}\sum_{j=0}^{H}C^{1,1}_{i,j}\left\|(i^{\prime},j^{\prime})-f_{1}\frac{(X^{1,1}_{i,j,0},X^{1,1}_{i,j,1})}{X^{1,1}_{i,j,2}}\right\|,\vspace{-1mm}$
@@
$R^{*},t^{*}=\operatorname*{arg\,min}_{\sigma,R,t}\sum_{i}C^{1,1}_{i}C^{1,2}_{i}\left\|\sigma(RX^{1,1}_{i}+t)-X^{1,2}_{i}\right\|^{2},\vspace{-2mm}$
@@ 3.4 Global Alignment
$\chi^{*}=\operatorname*{arg\,min}_{\chi,P,\sigma}\sum_{e\in\mathcal{E}}\sum_{v\in e}\sum_{i=1}^{HW}C^{v,e}_{i}\left\|\chi_{i}^{v}-\sigma_{e}P_{e}X^{v,e}_{i}\right\|.\vspace{-3mm}$ (5)
@@ 4 Experiments with DUSt3R
Training details. During each epoch, we randomly sample an equal number of pairs from each dataset to equalize disparities in dataset sizes. We wish to feed relatively high-resolution images to our network, say 512 pixels in the largest dimension. To mitigate the high cost associated with such input, we train our network sequentially, first on 224$\times$224 images and then on larger 512-pixel images. We randomly select the image aspect ratios for each batch (e.g. 16/9, 4/3, etc), so that at test time our network is familiar with different image shapes. We simply crop images to the desired aspect-ratio, and resize so that the largest dimension is 512 pixels.
@@ 4.2 Multi-view Pose Estimation
Baselines and metrics. We compare DUSt3R pose estimation results, obtained either from PnP-RANSAC or global alignment, against the learning-based RelPose [177], PoseReg [140] and PoseDiffusion [140], and structure-based PixSFM [59], COLMAP+SPSG (COLMAP [107] extended with SuperPoint [26] and SuperGlue [100]). Similar to [140], we report the Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) for each image pair to evaluate the relative pose error and select a threshold $\tau=15$ to report RTA@$15$ and RRA@$15$. Additionally, we calculate the mean Average Accuracy (mAA)@$30$, defined as the area under the curve accuracy of the angular differences at $min({\rm RRA}@30,{\rm RTA}@30)$.
@@
Monodepth2 [37] SS 23.91 75.22 11.42 86.90 56.49 35.18 16.19 74.50 31.20 47.42
SC-SfM-Learners [6] SS 16.92 77.28 11.83 86.61 21.11 71.40 13.79 79.57 22.29 64.30
SC-DepthV3 [121] SS 14.20 81.27 11.79 86.39 12.58 88.92 12.34 84.80 16.28 79.67
MonoViT[182] SS - - 09.92 90.01 - - - - -
@@
Table 2: Left: Monocular depth estimation on multiple benchmarks. D-Supervised, SS-Self-supervised, T-transfer (zero-shot). (Parentheses) refers to training on the same set. Right: Multi-view pose regression on the CO3Dv2 [94] and RealEst10K [186] with 10 random frames.
@@
Pose Range Intrinsics rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$ $\downarrow$ $\tau\uparrow$ $\downarrow$ $\tau\uparrow$ time (s)$\downarrow$
@@ 4.6 Ablations
Table 4: MVS results on the DTU dataset, in mm. Traditional handcrafted methods (a) have been overcome by learning-based approaches (b) that train on this specific domain.